# SMooGPT: Stylized Motion Generation using Large Language Models

### v1

@@ Abstract.
Stylized motion generation is actively studied in computer graphics, especially benefiting from the rapid advances in diffusion models. The goal of this task is to produce a novel motion respecting both the motion content and the desired motion style, e.g., “walking in a loop like a Monkey”. Existing research attempts to address this problem via motion style transfer or conditional motion generation. They typically embed the motion style into a latent space and guide the motion implicitly in a latent space as well. Despite the progress, their methods suffer from low interpretability and control, limited generalization to new styles, and fail to produce motions other than “walking” due to the strong bias in the public stylization dataset. In this paper, we propose to solve the stylized motion generation problem from a new perspective of reasoning-composition-generation, based on our observations: i) human motion can often be effectively described using natural language in a body-part centric manner, ii) LLMs exhibit a strong ability to understand and reason about human motion, and iii) human motion has an inherently compositional nature, facilitating the new motion content or style generation via effective recomposing. We thus propose utilizing body-part text space as an intermediate representation, and present SMooGPT, a fine-tuned LLM, acting as a reasoner, composer, and generator when generating the desired stylized motion. Our method executes in the body-part text space with much higher interpretability, enabling fine-grained motion control, effectively resolves potential conflicts between motion content and style, and generalizes well to new styles thanks to the open-vocabulary ability of LLMs. Comprehensive experiments and evaluations, and a user perceptual study, demonstrate the effectiveness of our approach, especially under the pure text-driven stylized motion generation.
@@ 1. Introduction
Human motion can generally be decomposed into two components: content and style. The content describes what the movement entails (e.g., walking, waving), whereas the style characterizes how the movement is executed, often reflecting personality traits (e.g., elderly, childlike) or emotional states (e.g., happy, angry). To automatically and efficiently generate stylized motion, existing works (Jang et al., 2022; Wen et al., 2021; Aberman et al., 2020) have primarily focused on transferring style from an input stylized motion sequence to a target content motion sequence. Recent advances in text-to-motion diffusion models (Tevet et al., 2023; Chen et al., 2023) enable direct stylized motion generation (Zhong et al., 2024; Sawdayee et al., 2025; Guo et al., 2025) using both motion sequences and texts as content and style inputs.
@@
In this paper, we propose solving this motion stylization problem from a new perspective of body part motion analysis. Our key insight is that both motion content and style can be interpreted as fine-grained control over body-part movements, and these 4D motions can be further intuitively expressed with natural language. For example, as shown in Fig. 2(a) and (b), the “Aeroplane” style is characterized primarily by the upward extension of both arms, while the “Chicken” motion can be interpreted with the same set of body parts in the text domain. Given the intrinsic correlation among motion, body parts, and natural language, we investigate a fundamental research question: can stylized motion generation be effectively achieved through body-part centric textual descriptions?
@@
At the same time, Large Language Models (LLMs) exhibit strong capabilities in understanding and reasoning about human poses (Lin et al., 2024; Feng et al., 2024) (see an example in Fig. 2(c)) — including stylized ones — and their representational scope extends well beyond that of existing motion style datasets (Mason et al., 2022a; Xia et al., 2015). Crucially, LLMs are particularly well-suited for decomposing and recomposing motion style and content at the text domain, enabling more interpretable and conflict-free compositional representations of stylization. Given these strengths, it is compelling to leverage the learned knowledge of human motion and reasoning abilities of LLMs to guide stylized motion generation in a more human-interpretable manner. On the other hand, interpreting stylized motion at the body-part level enables more effective use of existing datasets. While a new motion style may be unseen at the global level, many of its constituent body-part motions are likely to already exist within the dataset.
@@
Building upon these insights, we propose SMooGPT, a fine-tuned LLM, to solve the stylized motion generation problem following a novel reasoning-composition-generation methodology. The core of our method is to harness the reasoning capabilities of LLMs to translate the desired motion content and style into body-part texts, which are then used to generate the desired motion. Specifically, given the motion content and style represented as motion sequences or texts, SMooGPT is first instructed to reason out the corresponding body-part texts for each of them, acting as a motion reasoner. Next, instead of directly combining body-part texts from the content and style, which might potentially cause conflict, SMooGPT is further instructed to compose the two sets of body-part texts into unified and conflict-free body-part texts that respect both the content and the style, acting as a composer. Finally, given the composed body-part texts, SMooGPT is instructed to generate the expected stylized motion, acting as a motion generator. To enable the versatile abilities, we have fine-tuned an LLM with pre- and post-training stages on curated datasets derived from HumanML3D (Guo et al., 2022a) and 100STYLE (Mason et al., 2022a). The goal of pre-training is to bridge the text and motion modalities, enabling motion understanding and reasoning in text domain, while the post-training is to enable composition and generation. By leveraging the strong understanding and reasoning capabilities of LLMs within the body-part text domain, our approach effectively resolves conflicts between motion content and style, and demonstrates a certain degree of generalization to new styles and contents - challenges that remain difficult for state-of-the-art methods at present.
@@ 2. Related Work
Human motion synthesis focuses on generating diverse and realistic human-like movements. Traditional approaches rely on concatenation and retrieval techniques, including motion graphs (Kovar et al., 2023; Heck and Gleicher, 2007; Shin and Oh, 2006) and motion matching (Büttner and Clavet, 2015). Recently, learning-based methods have been widely adopted to generate high-quality and realistic motions, utilizing models such as GANs (Li et al., 2022; Ghosh et al., 2021), autoregressive GPTs (Guo et al., 2022b; Jiang et al., 2023; Zhang et al., 2023, 2024), generative masked Transfomers (Pinyoanuntapong et al., 2024; Guo et al., 2024a; Li et al., 2024c), and diffusion models (Tevet et al., 2023; Zhang et al., 2022a; Chen et al., 2023; Zhou et al., 2025).
@@
Motion stylization (Aberman et al., 2020; Kim et al., 2024; Tang et al., 2023; Wen et al., 2021; Mason et al., 2022b) is the process of creating stylized motion based on content and style specifications. Early methods mainly focus on motion style transfer (Song et al., 2024; Tao et al., 2022; Aberman et al., 2020), typically leveraging existing content and style sequences for recombination. Specifically, Aberman et al.  (2020) proposed the use of AdaIN to separate motion style from content and enable their flexible re-composition, which has since been widely adopted by many follow-up works (Guo et al., 2024b). Motion Puzzle (Jang et al., 2022) realizes a framework that can control the style of individual body parts. A major limitation of these models is their reliance on specialized style datasets (Mason et al., 2022b; Xia et al., 2015) with limited motion content, which restricts their applications
More recently, to extend the expressiveness of motion content, some approaches (Zhong et al., 2024; Guo et al., 2025; Li et al., 2024b; Sawdayee et al., 2025) leverage pre-trained text-to-motion models (Tevet et al., 2023; Chen et al., 2023) to generate stylized motions from text and style sequences. Specifically, SMooDi (Zhong et al., 2024) customizes a pre-trained text-to-motion model on the 100STYLE dataset to enable stylization across diverse motion content and styles. StyleMotif (Guo et al., 2025) and MulSMo (Li et al., 2024b) extend the style input to a multimodal setting based on SMooDi  [^1] Code from StyleMotif and MulSMo was not publicly available at the time of submission. However, these methods are still face challenges in generalizing to diverse motion styles that are unseen in the dataset. In contrast, we utilize the body-part textual domain and propose an reasoning-compositing-generation framework, which harnesses the strong reasoning ability of LLMs, ensuring the generalization of our approach to unseen styles in the wild.
@@
Large Language Models (Devlin et al., 2019; Dai et al., 2019; Raffel et al., 2020; Brown et al., 2020; Zhang et al., 2022b; Touvron et al., 2023) have shown remarkable capabilities in textual comprehension and reasoning. For instance, BERT (Devlin et al., 2019) adopts a bidirectional transformer structure and pre-trains the model applicable for various downstream tasks. T5 (Raffel et al., 2020) presents a unified solution by converting all language-related tasks into a text-to-text formulation. In a later stage, the pre-trained LLMs are adapted for multimodal tasks in various domains - image editing (Fu et al., 2023), layout generation (Littlefair et al., 2025), mathematical reasoning (Romera-Paredes et al., 2024), mesh generation (Wang et al., 2024), garment generation (Bian et al., 2024), and motion generation (Wu et al., 2024; Yu et al., 2025; Li et al., 2024a), to name a few.
More relevant to our scenario, Mojito (Shan et al., 2025) presents an interactive motion capture and behavioral analysis system with the assistance of LLMs. MotionGPT (Jiang et al., 2023; Zhang et al., 2024) treats human motion as a foreign language and leverages the language understanding and transfer abilities of pre-trained language models to support various tasks, like motion in-betweening, text-to-motion generation, and motion-to-text captioning. It also establishes the popular methodology of motion tokenization and LLM-based interaction. Based on this methodology, MotionChain (Jiang et al., 2024) supports multi-turn interactions by sampling single-turn data into multi-turn training examples, while Chen et al.  (2024) includes more modalities (e.g., emotion, audio, and hands) in the interaction with human motion in LLMs. Although our technical implementation is similar to MotionGPT (Jiang et al., 2023), an obvious distinction is that we introduce the body-part textual domain and take it as a common intermediate representation to achieve fine-grained content and style control in the stylized motion generation task. Besides, none of the above models can naturally handle this task, and our SMooGPT complements existing work.
@@ 3.1. Body-part Space
Formally, our body-part space is formulated as a triplet $\{\mathbf{m},\mathbf{T}_{g},\{\mathbf{T}_{p}\}\}$. As illustrated in Fig. 3, each motion $\mathbf{m}$ in body-part space can be interpreted as a composition of part-specific behaviors collectively expressing the full-body movement. The global text $\mathbf{T}_{g}$ primarily captures the overall semantics of the motion along with high-level stylistic modifiers (e.g., “walk forward and then walk backward in a flying pose”), while the body-part texts $\{\mathbf{T}_{p}\}$ provide detailed, localized descriptions for individual body parts. Importantly, motion style is manifested not only in static poses but also in dynamic movement patterns. Therefore, our body-part textual descriptions are designed to capture both the spatial configuration of each body part and its dynamic properties, such as movement direction and rhythm. To demonstrate the compositional nature of our body-part representation, Fig. 3 (bottom-left) presents an example in which modifying only a few body-part descriptions—e.g., changing the left arm motion from pointing “backward” to “forward”—produces a new stylized motion while preserving the underlying content. This example underscores the flexibility of our design: part-level representations improve the reusability of existing datasets and enable compositional generation, allowing novel motion styles to be synthesized by recombining previously observed body-part movements.
@@ 3.2. Generation Methodology
  • Generation: having the unified body-part texts, the last task is to simply translate them into the motion space, producing the resulting motion. Note that the body-part specific motion tokens will be decoded back to the motion space using the pre-trained decoders (see Sec. 4.1).
@@ 4.1. Motion Tokenization
where $\mathrm{sg}[\cdot]$ represents the stop-gradient operation, and $\beta$ is the weight for the commitment loss. This part-wise tokenization provides synchronized discrete token sequences across different body parts, facilitating fine-grained motion modeling and reasoning in the subsequent stages.
@@ 4.3. Post-training with Instruction Following
As shown in Fig. 4(a) and stated in Sec. 3.2, our key idea is to fully exploit the intermediate body-part text space for stylized motion generation. To this end, we design several instruction-following templates for the above bidirectional translations, as well as the extra global text to body-part texts predictions. In what follows, we only demonstrate one example, leaving detailed instruction templates in the supplementary.
Input: Global motion text $\mathbf{T}_{g}$>
With the aforementioned pre- and post-training on our style-agnostic dataset, our SMooGPT becomes a general-purpose motion reasoner and generator that operates on body-part texts. As for the composition ability, we design an additional instruction following task, where we fine-tune SMooGPT with content body-part texts $\mathbf{T}_{p}^{c}$ and style body-part texts $\mathbf{T}_{p}^{s}$ as shown below. Note that superscripts $c$ and $s$ differentiate the content text from the style text, and please refer to the supplementary to know their construction.
@@ 5. Results and Evaluation
With our SMooGPT, we have generated many appealing stylized motions, ranging from “Kungfu kick”, “Aeroplane jumping”, to “Lazy punch”, as presented in Figs. 1 and 5. Note that, styles like “Marching”, “Stage Performance”, “Cheering”, are not labeled or explicitly present in our HumanML3D training data, which reflects our method’s ability to generalize to unseen or underrepresented styles, based solely on their natural language descriptions, benefiting from LLM’s text reasoning ability. We also show simplified body-part texts in Fig. 5, and the full version can be found in the supplementary. Although our method is versatile in stylized motion generation by accepting both motion sequences and text descriptions as conditions, we specialize in pure text-guided motion stylization. Thus, we mainly demonstrate this application in the gallery (Fig. 5). Additionally, we compare baseline methods from motion-guided stylization in Sec. 5.1. Ablation studies are conducted in Sec. 5.2, validating our technical design. A user perceptual study (Sec. 5.3) also supports our superior performance. Some critical discussions are presented in Sec. 5.4. A supplemental video is provided for better visualization of the motion.
@@
We choose HumanML3D dataset (Guo et al., 2022a) as our training set for both the pre- and post-training. We utilize its consistent root-velocity motion representation and process the dataset to obtain our body-part space representation. For evaluation, we use the HumanML3D test set as the content source and randomly pair styles represented as texts or motion sequences from the test set of 100STYLE dataset (Mason et al., 2022a).
The detailed data construction process for both training and testing sets is in the supplementary.
Table 1. Quantitative Comparison Results. We compare with baselines on text-guided and motion-guided stylization. Motion content preservation, realism, and style reflection metrics are reported.
Method SRA $\uparrow$ MM Dist $\downarrow$ R-Precision (Top-3) $\uparrow$ FS-Ratio $\downarrow$
Text-Guided Stylization
ChatGPT + MotionGPT (Jiang et al., 2023) 7.61 4.625 0.591 0.124
ChatGPT + MLD (Chen et al., 2023) 4.82 4.313 0.605 0.131
SMooGPT (Ours) 33.33 5.096 0.563 0.141

### v2

@@ Abstract.
Stylized motion generation is actively studied in computer graphics, especially benefiting from the rapid advances in diffusion models. The goal of this task is to produce a novel motion respecting both the motion content and the desired motion style, e.g., “Walking in a loop like a Monkey”. Existing research attempts to address this problem via motion style transfer or conditional motion generation. They typically embed the motion style into a latent space and guide the motion implicitly in a latent space as well. Despite the progress, their methods suffer from low interpretability and control, limited generalization to new styles, and fail to produce motions other than “walking” due to the strong bias in the public stylization dataset. In this paper, we propose to solve the stylized motion generation problem from a new perspective of reasoning-composition-generation, based on our observations: i) human motion can often be effectively described using natural language in a body-part centric manner, ii) LLMs exhibit a strong ability to understand and reason about human motion, and iii) human motion has an inherently compositional nature, facilitating the new motion content or style generation via effective recomposing. We thus propose utilizing body-part text space as an intermediate representation, and present SMooGPT, a fine-tuned LLM, acting as a reasoner, composer, and generator when generating the desired stylized motion. Our method executes in the body-part text space with much higher interpretability, enabling fine-grained motion control, effectively resolving potential conflicts between motion content and style, and generalizes well to new styles thanks to the open-vocabulary ability of LLMs. Comprehensive experiments and evaluations, and a user perceptual study, demonstrate the effectiveness of our approach, especially under the pure text-driven stylized motion generation.
@@ 1. Introduction
Human motion can generally be decomposed into two components: content and style. The content describes what the movement entails (e.g., walking, waving), whereas the style characterizes how the movement is executed, often reflecting personality traits (e.g., elderly, childlike) or emotional states (e.g., happy, angry). To automatically and efficiently generate stylized motion, existing works (Jang et al., 2022; Wen et al., 2021; Aberman et al., 2020) have primarily focused on transferring style from an input stylized motion sequence to a target content motion sequence. Recent advances in text-to-motion diffusion models (Tevet et al., 2023; Chen et al., 2023) enable direct stylized motion generation (Zhong et al., 2024; Sawdayee et al., 2025; Guo et al., 2025; Li et al., 2024b) using both motion sequences and texts as content and style inputs.
@@
In this paper, we propose solving this motion stylization problem from a new perspective of body part motion analysis. Our key insight is that both motion content and style can be interpreted as fine-grained control over body-part movements, and these 3D motions can be further intuitively expressed with natural language. For example, as shown in Fig. 2(a) and (b), the “Aeroplane” style is characterized primarily by the upward extension of both arms, while the “Chicken” motion can be interpreted with the same set of body parts in the text domain. Given the intrinsic correlation among motion, body parts, and natural language, we investigate a fundamental research question: can stylized motion generation be effectively achieved through body-part centric textual descriptions?
@@
At the same time, Large Language Models (LLMs) exhibit strong capabilities in understanding and reasoning about human poses (Feng et al., 2024; Li et al., 2025a) and motions (Lin et al., 2024; Li et al., 2025b; Chen et al., 2025a) (see an example in Fig. 2(c))—including stylized ones—and their representational scope extends well beyond that of existing motion style datasets (Mason et al., 2022a; Xia et al., 2015). Crucially, recent work (Li et al., 2025c) empirically confirms that LLMs possess the capability to accurately deconstruct complex motion instructions into part-specific descriptions. Capitalizing on this unique advantage, it is compelling to leverage the learned motion knowledge and reasoning abilities of LLMs to decompose and recompose motion style and content within the text domain. This enables more interpretable and conflict-free compositional representations of stylization. On the other hand, interpreting stylized motion at the body-part level enables more effective use of existing datasets. While a new motion style may be unseen at the global level, many of its constituent body-part motions are likely to already exist within the dataset.
@@
Building upon these insights, we propose SMooGPT, a fine-tuned LLM, to solve the stylized motion generation problem following a novel reasoning-composition-generation methodology. The core of our method is to harness the reasoning capabilities of LLMs to translate the desired motion content and style into body-part texts, which are then used to generate the desired motion. Specifically, given the motion content and style represented as motion sequences or texts, SMooGPT is first instructed to reason out the corresponding body-part texts for each of them, acting as a motion reasoner. Next, instead of directly combining body-part texts from the content and style, which might potentially cause conflict, SMooGPT is further instructed to compose the two sets of body-part texts into unified and conflict-free body-part texts that respect both the content and the style, acting as a composer. Finally, given the composed body-part texts, SMooGPT is instructed to generate the expected stylized motion, acting as a motion generator. To enable the versatile abilities, we have fine-tuned an LLM with pre- and post-training stages on curated datasets derived from HumanML3D (Guo et al., 2022a) and 100STYLE (Mason et al., 2022a). The goal of pre-training is to bridge the text and motion modalities, enabling motion understanding and reasoning in the text domain, while the post-training is to enable composition and generation. Furthermore, a lightweight diffusion head is leveraged to resolve inconsistencies between independent body-part motions, ensuring a harmonized and coordinated final stylized motion. By leveraging the understanding and reasoning capabilities of LLMs within the body-part text domain, our approach effectively resolves conflicts between motion content and style, and demonstrates a certain degree of generalization to new styles and contents - challenges that remain difficult for state-of-the-art methods at present.
@@ 2. Related Work
Human motion synthesis focuses on generating diverse and realistic human-like movements. Traditional approaches rely on concatenation and retrieval techniques, including motion graphs (Kovar et al., 2023; Heck and Gleicher, 2007; Shin and Oh, 2006) and motion matching (Büttner and Clavet, 2015). Recently, learning-based methods have been widely adopted to generate high-quality and realistic motions, utilizing models such as GANs (Li et al., 2022; Ghosh et al., 2021), autoregressive GPTs (Guo et al., 2022b; Jiang et al., 2023; Zhang et al., 2023, 2024), generative masked Transfomers (Pinyoanuntapong et al., 2024; Guo et al., 2024a; Li et al., 2024c; Meng et al., 2025), and diffusion models (Tevet et al., 2023; Zhang et al., 2022a; Zhou et al., 2025; Zhong et al., 2025).
@@
Motion stylization (Aberman et al., 2020; Kim et al., 2024; Tang et al., 2023; Wen et al., 2021; Mason et al., 2022b; Wu et al., 2025; Chen et al., 2025b) is the process of creating stylized motion based on content and style specifications. Early methods mainly focus on motion style transfer (Song et al., 2024; Tao et al., 2022; Aberman et al., 2020), typically leveraging existing content and style sequences for recombination. Specifically, Aberman et al.  (2020) proposed the use of AdaIN to separate motion style from content and enable their flexible re-composition, which has since been widely adopted by many follow-up works (Guo et al., 2024b). Motion Puzzle (Jang et al., 2022) realizes a framework that can control the style of individual body parts. A major limitation of these models is their reliance on specialized style datasets (Mason et al., 2022b; Xia et al., 2015) with limited motion content, which restricts their applications
More recently, to extend the expressiveness of motion content, some approaches (Zhong et al., 2024; Guo et al., 2025; Li et al., 2024b; Sawdayee et al., 2025) leverage pre-trained text-to-motion models (Tevet et al., 2023; Chen et al., 2023) to generate stylized motions from text and style sequences. Specifically, SMooDi (Zhong et al., 2024) customizes a pre-trained text-to-motion model on the 100STYLE dataset to enable stylization across diverse motion content and styles. StyleMotif (Guo et al., 2025) and MulSMo (Li et al., 2024b) extend the style input to a multimodal setting based on SMooDi  [^1] Code from StyleMotif and MulSMo was not publicly available at the time of submission.. However, these methods still face challenges in generalizing to diverse motion styles that are unseen in the dataset. In contrast, we utilize the body-part textual domain and propose an reasoning-compositing-generation framework, which harnesses the reasoning ability of LLMs, ensuring the generalization of our approach to unseen styles in the wild.
@@
Large Language Models (Devlin et al., 2019; Dai et al., 2019; Raffel et al., 2020; Brown et al., 2020; Zhang et al., 2022b; Touvron et al., 2023) have shown remarkable capabilities in textual comprehension and reasoning. For instance, T5 (Raffel et al., 2020) presents a unified solution by converting all language-related tasks into a text-to-text formulation. In a later stage, the pre-trained LLMs are adapted for multimodal tasks in various domains - mesh generation (Wang et al., 2024), garment generation (Bian et al., 2024), and motion generation (Shan et al., 2025; Wu et al., 2024; Yu et al., 2025; Li et al., 2024a).
More relevant to our scenario, MotionGPT (Jiang et al., 2023; Zhang et al., 2024; Zhu et al., 2025) treats human motion as a foreign language and leverages the language understanding and transfer abilities of pre-trained language models to support various tasks, like motion in-betweening, text-to-motion generation, and motion-to-text captioning. It also establishes the popular methodology of motion tokenization and LLM-based interaction. Based on this methodology, MotionChain (Jiang et al., 2024) supports multi-turn interactions by sampling single-turn data into multi-turn training examples, while Chen et al.  (2024) includes more modalities (e.g., emotion, audio, and hands) in the interaction with human motion in LLMs. Although our technical implementation is similar to MotionGPT (Jiang et al., 2023), an obvious distinction is that we introduce the body-part textual domain and take it as a common intermediate representation to achieve fine-grained content and style control in the stylized motion generation task. Besides, none of the above models can naturally handle this task, and SMooGPT complements existing work.
@@ 3.1. Body-part Space
Formally, our body-part space is formulated as a triplet $\{\mathbf{m},\mathbf{T}_{g},\{\mathbf{T}_{p}\}\}$. As illustrated in Fig. 3, each motion $\mathbf{m}$ in the body-part space can be interpreted as a composition of part-specific behaviors collectively expressing the full-body movement. The global text $\mathbf{T}_{g}$ primarily captures the overall semantics of the motion along with high-level stylistic modifiers (e.g., “walk forward and then walk backward in a flying pose”), while the body-part texts $\{\mathbf{T}_{p}\}$ provide detailed, localized descriptions for individual body parts. Importantly, motion style is manifested not only in static poses but also in dynamic movement patterns. Therefore, our body-part textual descriptions are designed to capture both the spatial configuration of each body part and its dynamic properties, such as movement direction and rhythm. To demonstrate the compositional nature of our body-part representation, Fig. 3 (bottom-left) presents an example in which modifying only a few body-part descriptions—e.g., changing the left arm motion from pointing “backward” to “forward”—produces a new stylized motion while preserving the underlying content. This example underscores the flexibility of our design: part-level representations improve the reusability of existing datasets and enable compositional generation, allowing novel motion styles to be synthesized by recombining previously observed body-part movements.
@@ 3.2. Generation Methodology
  • Generation: having the unified body-part texts, the last task is to simply translate them into the motion space, producing the resulting motion. Note that the body-part specific motion tokens will first go through a diffusion denoising process (Sec. 4.5) to obtain the harmonized and coordinated motion tokens, which are further decoded back to the motion space using the pre-trained decoders (see Sec. 4.1).
@@ 4.1. Motion Tokenization
where $\mathrm{sg}[\cdot]$ represents the stop-gradient operator, and $\beta$ is the weight for the commitment loss. This part-wise tokenization provides synchronized discrete token sequences across different body parts, facilitating fine-grained motion modeling and reasoning in the subsequent stages.
@@ 4.3. Post-training with Instruction Following
As shown in Fig. 4(a) and stated in Sec. 3.2, our key idea is to fully exploit the intermediate body-part text space for stylized motion generation. To this end, we design several instruction-following templates for the above bidirectional translations, as well as the extra global text to body-part texts predictions
With the aforementioned pre- and post-training on our style-agnostic dataset, SMooGPT becomes a general-purpose motion reasoner and generator that operates on body-part texts. As for the composition ability, we design an additional instruction following task, where we fine-tune SMooGPT with content body-part texts $\mathbf{T}_{p}^{c}$ and style body-part texts $\mathbf{T}_{p}^{s}$ as shown below. Note that superscripts $c$ and $s$ differentiate the content text from the style text, and please refer to the supplementary to know their construction.
@@ 4.4. Language Model Training Details
4.5. Diffusion-based Motion Refinement
To bridge the gap between discrete LLM outputs and continuous motion dynamics, we introduce a lightweight diffusion head preceding the decoder. Specifically, the LLM-predicted body-part latents $\{c_{p}\}$ are aggregated into a unified condition vector $\bm{z}_{\text{cond}}$. We then employ a conditional denoiser $\epsilon_{\theta}(\bm{z}_{t},t,\bm{z}_{\text{cond}})$ to iteratively refine a latent vector starting from Gaussian noise $\bm{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$. This process yields a globally coordinated latent $\bm{z}_{0}$, which is subsequently decomposed back into body-part features and decoded by the VQ-decoder $\mathcal{D}$ to generate the final smooth motion sequence $\bm{m}$.
Figure 7. Visual Ablation. Qualitative comparison of ablated variants under the same content and style texts. Removing the body-part space (w/o BP), pre-training (w/o PT), or diffusion head (w/o DF) leads to degraded style expression, realism, or content–style consistency, while the full model produces coherent and realistic stylized motion. Please refer to the accompanying video for detailed comparisons.
@@ 5. Results and Evaluation
With our SMooGPT, we have generated many appealing stylized motions, ranging from “Kungfu kick”, “Aeroplane jumping”, to “Lazy punch”, as presented in Figs. 1 and 5. Note that, styles like “Marching”, “Stage Performance”, “Cheering”, are not labeled or explicitly present in our HumanML3D training data, which reflects our method’s ability to generalize to unseen or underrepresented styles, based solely on their natural language descriptions, benefiting from LLM’s text reasoning ability. We also show simplified body-part texts in Fig. 5, and the full version can be found in the supplementary. Although our method is versatile in stylized motion generation by accepting both motion sequences and text descriptions as conditions, we specialize in pure text-guided motion stylization. Thus, we mainly demonstrate this application in the gallery (Fig. 5). Additionally, we compare baseline methods from motion-guided stylization in Sec. 5.1. Ablation studies are conducted in Sec. 5.2, validating our technical design. A user perceptual study (Sec. 5.3) also supports our superior performance. Some critical discussions are presented in Sec. 5.4 and the supplementary. A supplemental video is provided for better visualization of the motion.
@@
We choose HumanML3D dataset (Guo et al., 2022a) as our training set for both the pre- and post-training. We utilize its consistent root-velocity motion representation and process the dataset to obtain our body-part space representation. For evaluation, we use the HumanML3D test set as the content source and randomly pair styles represented as texts or motion sequences from the test set of 100STYLE dataset (Mason et al., 2022a). The detailed data construction process is in the supplementary.
Table 1. Quantitative Comparison Results. We compare with baselines on text-guided stylization. Motion content preservation ($M$ Dist $\downarrow$, R-Precision (Top-3) $\uparrow$, CLIP-scoreea $\uparrow$), realism (FS-Ratio $\downarrow$), and style reflection (SRA-L $\uparrow$) metrics are reported. The best results are highlighted in bold.
Method SRA-L $\uparrow$ MM Dist $\downarrow$ R-Precision (Top-3) $\uparrow$ CLIP-score $\uparrow$ FS-Ratio $\downarrow$
ChatGPT + MotionGPT (Jiang et al., 2023) 11.562 1.225 0.373 0.614 0.0984
ChatGPT + ParCo (Zou et al., 2024) 13.359 1.212 0.405 0.622 0.111
ChatGPT + MLD (Chen et al., 2023) 9.609 1.246 0.289 0.652 0.181
SMooGPT (Ours) 17.344 1.132 0.669 0.670 0.0832
Evaluation metrics.
We evaluate our method along three dimensions: content preservation, style reflection, and realism. For the assessment of content preservation, we employ metrics used in (Chen et al., 2023): motion-retrieval precision (R-Precision), Multi-modal Distance (MM Dist). Moreover, we utilize the CLIP-score (Meng et al., 2025) to evaluate the alignment between the text content and the stylized motion. Additionally, we exploit the foot skating ratio (FS-Ratio) proposed by (Karunratanakul et al., 2023) for our motion realism evaluation. For style reflection, we employ Style Recognition Accuracy (SRA) (Jang et al., 2022). Since the style classifier is trained on the 100STYLE dataset—which primarily consists of locomotion—it may be unreliable for non-locomotion content. Consequently, we calculate SRA by filtering for locomotion-related sequences within the HumanML3D test set, a metric we denote as SRA-L. We do not include the common FID metric for evaluating content preservation due to its inherent ambiguity in our task. See supplementary for the detailed discussion.
5.1. Comparison
Text-guided Stylization.
This task generates stylized motion based on both the content and style texts. All methods were trained only on the HumanML3D dataset, where styles are not labeled or explicitly present. Two concurrent works, StyleMotif (Guo et al., 2025) and MulSMo (Li et al., 2024b), are capable of producing stylized motion in this setting. However, both methods rely on style motion sequences from the 100STYLE dataset to align style labels. Furthermore, their source code was not publicly available at the time of submission. We thus create three baseline methods by integrating a pre-trained LLM (i.e., GPT-3.5 (OpenAI, 2024)) with three representative motion generators (i.e., MotionGPT (Jiang et al., 2023), ParCo (Zou et al., 2024), and MLD (Chen et al., 2023)). Strategically, we utilize content text from the HumanML3D dataset and style labels from the 100STYLE dataset. We first ask ChatGPT to combine the content and style label, and employ the respective motion generator to produce the corresponding motion according to the composed text. Statistical results are displayed in Tab. 1.
In terms of style reflection (SRA-L), SMooGPT outperforms the strongest baseline by approximately 29.8%. Furthermore, our method achieves superior performance across all motion content preservation and realism metrics. Specifically, we improve content consistency by reducing MM-Dist by 6.6% and increasing R-Precision by 65.2%, while also achieving the best motion realism with the lowest FS-Ratio (0.0832) among all compared methods. The corresponding visual results are presented in Fig. 6(top). In three typical cases, baseline methods either cannot preserve the content or fail to reflect the style.
Table 2. Quantitative Comparison on motion-guided stylization. Motion content preservation ($M$ Dist $\downarrow$, R-Precision (Top-3) $\uparrow$, CLIP-scoreea $\uparrow$), realism (FS-Ratio $\downarrow$), and style reflection (SRA-L $\uparrow$) metrics are reported. The best results are highlighted in bold.
Method SRA-L $\uparrow$ MM Dist $\downarrow$ R-Precision (Top-3) $\uparrow$ CLIP-score $\uparrow$ FS-Ratio $\downarrow$
Setting 1: Zero-Shot (Cross-Dataset)
SMooDi (Zhong et al., 2024) 7.110 1.185 0.504 0.640 0.112
SMooGPT (Ours) 12.695 1.153 0.481 0.659 0.104
Setting 2: Held-Out (Intra-Dataset)
SMooDi (Zhong et al., 2024) 14.775 1.157 0.509 0.651 0.118
SMooGPT (Ours) 15.527 1.150 0.474 0.660 0.110